what are the slopes of these lines, and the slopes of lines parallel to them?:
1.) y = 2/3x - 7
2.) y - 4 = -8/17(x-3)
3.) x = 4y + 2
4.) the line through (2, 8) (-4, -7)

Answer by Maths68(1474)

what are the slopes of these lines, and the slopes of lines parallel to them?:
=============================================
1.) y = 2/3x - 7
Compare the above equatoin with the Equation of the Stright line Standard form.
y=mx+b
Slope = m = 2/3
===============================================
2.) y - 4 = -8/17(x-3)
Rearrage the above euqation
y=-8/17(x-3)+4
y=-8/17(x)-(-8/17)(3)+4
y=-8/17(x)+(24/17)+4
y=-8/17(x)+(24+68)/17
y=-8/17(x)+92/17
Compare the above equation with the Equation of the Stright line Standard form.
y=mx+b
Slope=m=-8/17
============================================
3.) x = 4y + 2
Rearrage the equation
x-2=4y
(x-2)/4=4y/4
(x-2)/4=y
x/4-2/4=y
1/4(x)-1/2=y
y=1/4(x)-1/2
Compare the above equation with the Equation of the Stright line Standard form.
y=mx+b
Slope = m = 1/4
===========================================================================
4.) the line through (2, 8) (-4, -7)
Use equation of straight line point-slope form
m=y2-y1/x2-x1
m=-7-8/-4-2
m=-15/-6
m=5/2
Slope = m = 5/2
Since parallel lines have same slope, therefore lines parallel to them will have same slope.